# Quantum algorithms for viscosity solutions to nonlinear Hamilton-Jacobi equations based on an entropy penalisation method

### v1

@@ IV.1.2 Digital protocol to measure value at a point
The minimum precision in estimating $S(t,x_{a})$ when using $|u(t)\rangle$ constructed from $d\log N_{x}$ qubits and the normalisation constant of $S_{\nu}(t,x_{a})$ is $\epsilon_{\min}$, where a large enough $N_{x}$ is chosen so $1/N_{x}=O\left(\sqrt{\epsilon_{\min}/d}\right)$. Then, using Algorithm 6, the number of preparations of state $|v(t)\rangle$ in Step (2) of Algorithm 4 required to estimate $S(t,x_{a})$ to precision $\epsilon$ with success probability $1-\delta$ is $O(\ln(1/\delta)\nu^{2}/(\epsilon-\epsilon_{\min})^{2})$, and it is sufficient to choose the constant $\nu$ Algorithm 4 to be $O(\epsilon^{2}_{\min}/d^{2})$.
@@ IV.2.2 Digital protocol for gradient estimation
Let the minimum error to recover $\partial S(t,x_{a})/\partial x_{k}$ be $\epsilon_{\min}$. To estimate $\partial S(t,x_{a})/\partial x_{k}$ to precision $\epsilon$ requires $O(d\ln(1/\delta)/(\kappa^{2}(\epsilon-\epsilon_{\min})^{2}))$ copies of $|u(t)\rangle$ from Algorithm 4, and the same uses of the qubit-control gate $\exp(2i\epsilon\nu\hat{P}_{k}\otimes\sigma_{x})$, where $\hat{P_{k}}$ acts on a system of $O(d\log(N_{x}))$ qubits. In Algorithm 4 it is sufficient to choose $\nu=O(\epsilon^{2}_{\min}/d^{2})$ where $\epsilon_{\min}=d^{3/2}/N_{x}$.
@@ IV.3.1 Analog protocol to estimate $S_{\min}$
$|S_{\nu,\min}+\nu\log\|\mathbf{u}(t)\|^{2}|\lesssim\tilde{O}(\nu d),$ (62)
@@
$|S_{\min}+v\log\|\mathbf{u}(t)\|^{2}|\leq|S_{\min}-S_{\nu,\min}|+|S_{\nu,\min}+\nu\log\|\mathbf{u}(t)\|^{2}|\lesssim O(\tilde{\epsilon})+\tilde{O}(\nu d)=\epsilon_{\min},$ (63)
@@
$O\left(\frac{\nu^{2}}{\|\mathbf{u}(t)\|^{4}}\log\left(\frac{1}{\delta}\right)\frac{1}{(\epsilon_{S^{*}}-\epsilon_{\min})^{2}}\right),$ (64)
@@
$|S_{\min}+\nu\log\|\tilde{\mathbf{u}}(t)\|^{2}|\leq|S_{\min}-S_{\nu,\min}|+|S_{\nu,\min}+\nu\log\|\mathbf{u}(t)\|^{2}|+|\nu\log\|\mathbf{u}(t)\|^{2}-\nu\log\|\tilde{\mathbf{u}}(t)\|^{2}|$
$\leq O(\tilde{\epsilon})+\nu d\log(2\pi\nu/\det H_{S_{\nu}}(x^{*}))/2+\tilde{O}_{hot}(\nu)+\zeta\lesssim O(\tilde{\epsilon})+\tilde{O}(\nu d)+\zeta=\epsilon_{\min}+\zeta=\epsilon_{S^{*}},$ (65)
@@
One can relate $\zeta$ to the sampling cost in the following way. Let $\epsilon_{norm}$ denote the precision to which one can measure the normalisation $\|\mathbf{u}(t)\|^{2}$. Let $1-\delta$ be the success probability, then the total measurement cost (which is equivalent to the number of repetitions of the quantum simulation with respect to $\mathbf{H}(t)$, since the normalisation is estimated as a byproduct – it is the probability of measuring the final ancilla state in $\xi>0$) in the protocol in Fig. 1 is $O(\log(1/\delta)/\epsilon^{2}_{norm})$. Then we can use
$|\nu\log\|\mathbf{u}(t)\|^{2}-\nu\log\|\tilde{\mathbf{u}}(t)\|^{2}|\leq\zeta\implies\zeta\sim\nu\epsilon_{norm}/\|\mathbf{u}(t)\|^{2},$ (66)
so the sampling cost $O(\log(1/\delta)/\epsilon^{2}_{norm})\sim O(\nu^{2}\log(1/\delta)/(\zeta^{2}\|\mathbf{u}(t)\|^{4})$. From Eq. (IV.3.1) it is sufficient to choose $\nu=O(\epsilon^{2}_{\min}/d^{2})$ and $\zeta=\epsilon_{S^{*}}-\epsilon_{\min}(\nu)$, so the quantum sampling cost $O(\nu^{2}\log(1/\delta)/(\zeta^{2}\|\mathbf{u}(t)\|^{4})$ can be rewritten in terms of $\epsilon_{\min}(\nu)$ and $\epsilon_{S^{*}}$. ∎
@@ IV.3.2 Digital protocol to estimate $S_{\min}$
One can use a similar bound to Eq. (62) to estimate $S_{\nu,\min}$, except we can derive it in a much simpler way (similar to the LogSumExp (LSE) function $\log\sum_{j}\exp(y_{j})$, used widely in machine learning to construct a smooth approximation to a maximum function). It is simple to see the following (using $S_{\nu,\Delta,\min}(t)<S_{\nu,\Delta}(t,x_{j})$ for any $x_{j}$ that is not the minimum point) and Lemma 3 and Lemma 7:
$\bigg|\nu\ln\left(\|\mathbf{u}(t)\|_{DV}^{2}\right)+S_{\min}\bigg|\leq\bigg|\nu\ln\left(\|\mathbf{u}(t)\|_{DV}^{2}\right)+S_{\nu,\Delta,\min}\bigg|+\bigg|S_{\nu,\Delta,\min}-S_{\min}\bigg|\leq\bigg|\nu\log\left(\sum_{j}^{N_{x}^{d}}e^{-\frac{S_{\nu,\Delta}(t,x_{j})}{\nu}}\right)+S_{\nu,\Delta,\min}\bigg|+\tilde{\epsilon}$
$=\bigg|\nu\log\left(e^{-S_{\nu,\Delta,\min}(t)/\nu}\sum_{j}^{N_{x}^{d}}e^{\frac{-S_{\nu,\Delta}(t,x_{j})+S_{\nu,\Delta,\min}(t)}{\nu}}\right)+S_{\nu,\Delta,\min}\bigg|+\tilde{\epsilon}=\nu\log\left(\sum_{j}^{N_{x}^{d}}e^{\frac{-(S_{\nu,\Delta}(t,x_{j})-S_{\nu,\Delta,\min}(t))}{\nu}}\right)+\tilde{\epsilon}$
$\leq\nu\log(N_{x}^{d})+\tilde{\epsilon}=\nu d\log N_{x}+\tilde{\epsilon}=\epsilon_{\min},\qquad\nu=O(\tilde{\epsilon}^{2}/d^{2}),\qquad 1/N_{x}=O(\sqrt{\tilde{\epsilon}/d}),$ (68)
@@
Let minimum precision in estimating $S_{\min}$ using the normalisation constant from $S_{\nu}$ be $\epsilon_{\min}$ when one uses $\log N_{x}$ qubits to represent $|u(t)\rangle_{DV}$. Let $\epsilon_{S^{*}}$ be the total error in estimating $S_{\min}$ when we are using Algorithm 4 to estimate the normalisation constant $\|\mathbf{u}(t)\|_{DV}^{2}$. The number of preparations of state $|v(t)\rangle$ in Step (2) of Algorithm 2 to estimate the normalisation constant is
$O\left(\frac{\nu^{2}}{\|\mathbf{u}(t)\|^{4}}\log\left(\frac{1}{\delta}\right)\frac{1}{(\epsilon_{S^{*}}-\epsilon_{\min})^{2}}\right),$ (69)
where the constant $\nu$ used in the quantum simulation algorithm can be chosen $\nu=1/N_{x}^{4}$ and the resulting minimum error is $\epsilon_{\min}=\sqrt{\nu}d+\nu d\log(1/\nu)/4$.
@@
Here Algorithm 4 can be used to estimate the normalisation constant. The proof then proceeds in the same way as Lemma 17, except now $\epsilon_{min}$ is slightly modified, see Eq. (IV.3.2). Using Eq. (IV.3.2), one can for example choose $N_{x}^{2}=d/\tilde{\epsilon}$ and $\nu=\tilde{\epsilon}^{2}/d^{2}$, which implies $\nu=1/N_{x}^{4}$. Then the minimum error can be rewritten $\epsilon_{\min}=\sqrt{\nu}d+\nu d\log(1/\nu)/4=d/N_{x}^{2}+d\log(N_{x})/N_{x}^{4}\sim d/N_{x}^{2}$. ∎
@@ V Summary and outlook
In this work, we have established a unified framework for quantum simulation of viscosity solutions to nonlinear Hamilton–Jacobi equations with convex Hamiltonians, by combining the entropy penalisation method of Gomes and Valdinoci with quantum simulation of linear parabolic dynamics. This framework generalises the classical Cole–Hopf transformation, enabling a linear approximation of nonlinear viscous Hamilton–Jacobi equations that is valid for more general convex nonlinearities (quadratic and beyond) and arbitrary evolution times. This construction provides a relatively rare example of a nonlinear PDE class that admits an exact linear quantum representation, extending beyond perturbative truncations such as Carleman embeddings or other linear approximation methods that are not valid globally in time and strong nonlinearities.
@@ Appendix B First-order quantum simulation algorithms with sparse-access
which acts on a system of $\log N_{\eta}+d\log N_{x}$ qubits. Here $\hat{P}_{j}$ is a sparse shift matrix and $\hat{X}=\text{diag}(-1/2N_{x},\cdots,1/2N_{x})$ for the one-dimensional case, and $\hat{D}=\text{diag}(-1/2N_{\eta},\cdots,1/2N_{\eta})$.
@@
which acts on a system of $\log N_{\eta}+d\log N_{x}$ qubits. Here $\hat{P}_{j}$ is a sparse shift matrix and $\hat{X}=\text{diag}(-1/2N_{x},\cdots,1/2N_{x})$ for the one-dimensional case, and $\hat{D}=\text{diag}(-1/2N_{\eta},\cdots,1/2N_{\eta})$. The optimal cost for the simulation of $\exp(i\mathbf{H}t)$ for time-independent Hamiltonian simulation is $O(s_{H}\|\mathbf{H}\|_{\max}t)$, where $s_{H}$ is the row-sparsity of $\mathbf{H}$ and $\|\mathbf{H}\|_{\max}$ is the size of the maximum entry in $\mathbf{H}$. In this case, $s(\sum_{i}\mu_{i}\hat{P}_{i})=O(d)$, $s(\sum_{ij}\nu_{ij}\hat{P}_{i}\hat{P}_{j}\otimes\hat{D})=O(d^{2})$. The max-norm $\|\mathbf{H}\|_{max}\leq\|\sum_{i}\mu_{i}\hat{P}_{i}\otimes\mathbf{I}\|_{max}+\|\sum_{ij}\nu_{ij}\hat{P}_{i}\hat{P}_{j}\otimes\hat{D}\|_{max}$, where $\|\sum_{i}\mu_{i}\hat{P}_{i}\otimes\mathbf{I}\|_{max}\leq O(d|\mu^{max}|N_{x})$, $\|\sum_{ij}\nu_{ij}\hat{P}_{i}\hat{P}_{j}\otimes\hat{D}\|_{max}\leq O(\nu^{max}d^{2}N_{x}^{2}N_{\eta})$, with $\mu^{max}=\max_{i}\mu_{i}$, $\nu^{max}=\max_{ij}\nu_{ij}$. Thus, the optimal cost scales as $O(d^{3}N_{x}t(|\mu^{max}|+\nu^{max}dN_{x}N_{\eta}))$. Finally, to recover the state $|u(t)\rangle$ after the Hamiltonian simulation requires a post-selection process, which has the probability of success that can be amplified to $\sim\|\mathbf{u}(t)\|$.

### v2

@@ IV.1.2 Digital protocol to measure value at a point
The minimum precision in estimating $S(t,x_{a})$ when using $|u(t)\rangle$ constructed from $d\ln N_{x}$ qubits and the normalisation constant of $S_{\nu}(t,x_{a})$ is $\epsilon_{\min}$, where a large enough $N_{x}$ is chosen so $1/N_{x}=O\left(\sqrt{\epsilon_{\min}/d}\right)$. Then, using Algorithm 6, the number of preparations of state $|v(t)\rangle$ in Step (2) of Algorithm 4 required to estimate $S(t,x_{a})$ to precision $\epsilon$ with success probability $1-\delta$ is $O(\ln(1/\delta)\nu^{2}/(\epsilon-\epsilon_{\min})^{2})$, and it is sufficient to choose the constant $\nu$ Algorithm 4 to be $O(\epsilon^{2}_{\min}/d^{2})$.
@@ IV.2.2 Digital protocol for gradient estimation
Let the minimum error to recover $\partial S(t,x_{a})/\partial x_{k}$ be $\epsilon_{\min}$. To estimate $\partial S(t,x_{a})/\partial x_{k}$ to precision $\epsilon$ requires $O(d\ln(1/\delta)/(\kappa^{2}(\epsilon-\epsilon_{\min})^{2}))$ copies of $|u(t)\rangle$ from Algorithm 4, and the same uses of the qubit-control gate $\exp(2i\epsilon\nu\hat{P}_{k}\otimes\sigma_{x})$, where $\hat{P_{k}}$ acts on a system of $O(d\ln(N_{x}))$ qubits. In Algorithm 4 it is sufficient to choose $\nu=O(\epsilon^{2}_{\min}/d^{2})$ where $\epsilon_{\min}=d^{3/2}/N_{x}$.
@@ IV.3.1 Analog protocol to estimate $S_{\min}$
$|S_{\nu,\min}+\nu\ln\|\mathbf{u}(t)\|^{2}|\lesssim\tilde{O}(\nu d),$ (63)
@@
$|S_{\min}+v\ln\|\mathbf{u}(t)\|^{2}|\leq|S_{\min}-S_{\nu,\min}|+|S_{\nu,\min}+\nu\ln\|\mathbf{u}(t)\|^{2}|\lesssim O(\tilde{\epsilon})+\tilde{O}(\nu d)=\epsilon_{\min},$ (64)
@@
$O\left(\frac{\nu^{2}}{\|\mathbf{u}(t)\|^{4}}\ln\left(\frac{1}{\delta}\right)\frac{1}{(\epsilon_{S^{*}}-\epsilon_{\min})^{2}}\right),$ (65)
@@
$|S_{\min}+\nu\ln\|\tilde{\mathbf{u}}(t)\|^{2}|\leq|S_{\min}-S_{\nu,\min}|+|S_{\nu,\min}+\nu\ln\|\mathbf{u}(t)\|^{2}|+|\nu\ln\|\mathbf{u}(t)\|^{2}-\nu\ln\|\tilde{\mathbf{u}}(t)\|^{2}|$
$\leq O(\tilde{\epsilon})+\nu d\ln(2\pi\nu/\det H_{S_{\nu}}(x^{*}))/2+\tilde{O}_{hot}(\nu)+\zeta\lesssim O(\tilde{\epsilon})+\tilde{O}(\nu d)+\zeta=\epsilon_{\min}+\zeta=\epsilon_{S^{*}},$ (66)
@@
One can relate $\zeta$ to the sampling cost in the following way. Let $\epsilon_{norm}$ denote the precision to which one can measure the normalisation $\|\mathbf{u}(t)\|^{2}$. Let $1-\delta$ be the success probability, then the total measurement cost (which is equivalent to the number of repetitions of the quantum simulation with respect to $\mathbf{H}(t)$, since the normalisation is estimated as a byproduct – it is the probability of measuring the final ancilla state in $\xi>0$) in the protocol in Fig. 1 is $O(\ln(1/\delta)/\epsilon^{2}_{norm})$. Then we can use
$|\nu\ln\|\mathbf{u}(t)\|^{2}-\nu\ln\|\tilde{\mathbf{u}}(t)\|^{2}|\leq\zeta\implies\zeta\sim\nu\epsilon_{norm}/\|\mathbf{u}(t)\|^{2},$ (67)
so the sampling cost $O(\ln(1/\delta)/\epsilon^{2}_{norm})\sim O(\nu^{2}\ln(1/\delta)/(\zeta^{2}\|\mathbf{u}(t)\|^{4})$. From Eq. (IV.3.1) it is sufficient to choose $\nu=O(\epsilon^{2}_{\min}/d^{2})$ and $\zeta=\epsilon_{S^{*}}-\epsilon_{\min}(\nu)$, so the quantum sampling cost $O(\nu^{2}\ln(1/\delta)/(\zeta^{2}\|\mathbf{u}(t)\|^{4})$ can be rewritten in terms of $\epsilon_{\min}(\nu)$ and $\epsilon_{S^{*}}$. ∎
@@ IV.3.2 Digital protocol to estimate $S_{\min}$
One can use a similar bound to Eq. (63) to estimate $S_{\nu,\min}$, except we can derive it in a much simpler way (similar to the LogSumExp (LSE) function $\ln\sum_{j}\exp(y_{j})$, used widely in machine learning to construct a smooth approximation to a maximum function). It is simple to see the following (using $S_{\nu,\Delta,\min}(t)<S_{\nu,\Delta}(t,x_{j})$ for any $x_{j}$ that is not the minimum point) and Lemma 3 and Lemma 7:
$\bigg|\nu\ln\left(\|\mathbf{u}(t)\|_{DV}^{2}\right)+S_{\min}\bigg|\leq\bigg|\nu\ln\left(\|\mathbf{u}(t)\|_{DV}^{2}\right)+S_{\nu,\Delta,\min}\bigg|+\bigg|S_{\nu,\Delta,\min}-S_{\min}\bigg|\leq\bigg|\nu\ln\left(\sum_{j}^{N_{x}^{d}}e^{-\frac{S_{\nu,\Delta}(t,x_{j})}{\nu}}\right)+S_{\nu,\Delta,\min}\bigg|+\tilde{\epsilon}$
$=\bigg|\nu\ln\left(e^{-S_{\nu,\Delta,\min}(t)/\nu}\sum_{j}^{N_{x}^{d}}e^{\frac{-S_{\nu,\Delta}(t,x_{j})+S_{\nu,\Delta,\min}(t)}{\nu}}\right)+S_{\nu,\Delta,\min}\bigg|+\tilde{\epsilon}=\nu\ln\left(\sum_{j}^{N_{x}^{d}}e^{\frac{-(S_{\nu,\Delta}(t,x_{j})-S_{\nu,\Delta,\min}(t))}{\nu}}\right)+\tilde{\epsilon}$
$\leq\nu\ln(N_{x}^{d})+\tilde{\epsilon}=\nu d\ln N_{x}+\tilde{\epsilon}=\epsilon_{\min},\qquad\nu=O(\tilde{\epsilon}^{2}/d^{2}),\qquad 1/N_{x}=O(\sqrt{\tilde{\epsilon}/d}),$ (69)
@@
Let minimum precision in estimating $S_{\min}$ using the normalisation constant from $S_{\nu}$ be $\epsilon_{\min}$ when one uses $\ln N_{x}$ qubits to represent $|u(t)\rangle_{DV}$. Let $\epsilon_{S^{*}}$ be the total error in estimating $S_{\min}$ when we are using Algorithm 4 to estimate the normalisation constant $\|\mathbf{u}(t)\|_{DV}^{2}$. The number of preparations of state $|v(t)\rangle$ in Step (2) of Algorithm 2 to estimate the normalisation constant is
$O\left(\frac{\nu^{2}}{\|\mathbf{u}(t)\|^{4}}\ln\left(\frac{1}{\delta}\right)\frac{1}{(\epsilon_{S^{*}}-\epsilon_{\min})^{2}}\right),$ (70)
where the constant $\nu$ used in the quantum simulation algorithm can be chosen $\nu=1/N_{x}^{4}$ and the resulting minimum error is $\epsilon_{\min}=\sqrt{\nu}d+\nu d\ln(1/\nu)/4$.
@@
Here Algorithm 4 can be used to estimate the normalisation constant. The proof then proceeds in the same way as Lemma 17, except now $\epsilon_{min}$ is slightly modified, see Eq. (IV.3.2). Using Eq. (IV.3.2), one can for example choose $N_{x}^{2}=d/\tilde{\epsilon}$ and $\nu=\tilde{\epsilon}^{2}/d^{2}$, which implies $\nu=1/N_{x}^{4}$. Then the minimum error can be rewritten $\epsilon_{\min}=\sqrt{\nu}d+\nu d\ln(1/\nu)/4=d/N_{x}^{2}+d\ln(N_{x})/N_{x}^{4}\sim d/N_{x}^{2}$. ∎
@@ V Summary and outlook
In this work, we have established a unified framework for quantum simulation of viscosity solutions to nonlinear Hamilton–Jacobi equations with convex Hamiltonians, by combining the entropy penalisation method of Gomes and Valdinoci with quantum simulation of linear parabolic dynamics. This framework generalises the classical Cole–Hopf transformation, enabling a linear approximation of nonlinear viscous Hamilton–Jacobi equations that is valid for more general convex nonlinearities (quadratic and beyond) and arbitrary evolution times. This construction provides a relatively rare example of a nonlinear PDE class, with strong nonlinearity, small dissipation and admitting a global-in-time linear quantum formulation. This extends beyond perturbative truncations such as Carleman embeddings or other linear approximation methods that are in general not valid globally-in-time, with strong nonlinearities and small dissipation.
@@ Appendix B First-order quantum simulation algorithms with sparse-access
which acts on a system of $\ln N_{\eta}+d\ln N_{x}$ qubits. Here $\hat{P}_{j}$ is a sparse shift matrix and $\hat{X}=\text{diag}(-1/2N_{x},\cdots,1/2N_{x})$ for the one-dimensional case, and $\hat{D}=\text{diag}(-1/2N_{\eta},\cdots,1/2N_{\eta})$.
@@
which acts on a system of $\ln N_{\eta}+d\ln N_{x}$ qubits. Here $\hat{P}_{j}$ is a sparse shift matrix and $\hat{X}=\text{diag}(-1/2N_{x},\cdots,1/2N_{x})$ for the one-dimensional case, and $\hat{D}=\text{diag}(-1/2N_{\eta},\cdots,1/2N_{\eta})$. The optimal cost for the simulation of $\exp(i\mathbf{H}t)$ for time-independent Hamiltonian simulation is $O(s_{H}\|\mathbf{H}\|_{\max}t)$, where $s_{H}$ is the row-sparsity of $\mathbf{H}$ and $\|\mathbf{H}\|_{\max}$ is the size of the maximum entry in $\mathbf{H}$. In this case, $s(\sum_{i}\mu_{i}\hat{P}_{i})=O(d)$, $s(\sum_{ij}\nu_{ij}\hat{P}_{i}\hat{P}_{j}\otimes\hat{D})=O(d^{2})$. The max-norm $\|\mathbf{H}\|_{max}\leq\|\sum_{i}\mu_{i}\hat{P}_{i}\otimes\mathbf{I}\|_{max}+\|\sum_{ij}\nu_{ij}\hat{P}_{i}\hat{P}_{j}\otimes\hat{D}\|_{max}$, where $\|\sum_{i}\mu_{i}\hat{P}_{i}\otimes\mathbf{I}\|_{max}\leq O(d|\mu^{max}|N_{x})$, $\|\sum_{ij}\nu_{ij}\hat{P}_{i}\hat{P}_{j}\otimes\hat{D}\|_{max}\leq O(\nu^{max}d^{2}N_{x}^{2}N_{\eta})$, with $\mu^{max}=\max_{i}\mu_{i}$, $\nu^{max}=\max_{ij}\nu_{ij}$. Thus, the optimal cost scales as $O(d^{3}N_{x}t(|\mu^{max}|+\nu^{max}dN_{x}N_{\eta}))$. Finally, to recover the state $|u(t)\rangle$ after the Hamiltonian simulation requires a post-selection process, which has the probability of success that can be amplified to $\sim\|\mathbf{u}(t)\|$.